Find the perimeter of a rectangle, given that the length is z yards and the width is 3 yards shorter.
Answer by atif.muhammad(135)

Length = z
Width is 3 yards shorter = (z-3)
Perimenter = Length + Width + Length + Width = z + (z-3) + z +(z-3) = (4z-6) yards